# Simulating Molecular Single Vibronic Level Fluorescence Spectra with ab initio Hagedorn Wavepacket Dynamics

### v4

@@ 5 Conclusions
Because the vibrational Hagedorn dynamics in the global harmonic model is exact, it is easier to identify and isolate errors. Our analysis of the results with different harmonic approximations demonstrates the importance of including both mode distortion and Duschinsky rotation in evaluating molecular vibronic spectra, which serve as a sensitive probe for molecular structure and dynamics. In the case of anthracene, our results suggest that fine-tuning of mode mixing may be necessary to more accurately reproduce spectral features involving combination bands. In this sense, computing vibronic spectra could be used to assess and improve the accuracy of electronic structure methods.
The Hagedorn approach can be readily combined with the local harmonic approximation to capture mild anharmonic effects, though this requires a much more expensive propagation with on-the-fly electronic structure evaluations along the trajectory.61 The results from global harmonic models can indicate of when anharmonicity becomes significant and guide the selection of an appropriate ab initio method for further local harmonic studies.
Even within the global harmonic framework, an extension of our method could be useful for other experimental techniques involving vibrationally excited states, such as vibrationally promoted electronic resonance (VIPER) experiments,83, 84, 85 time-resolved photoelectron spectroscopy,86 and fluorescence-encoded infrared (FEIR) spectroscopy.87 More generally, Hagedorn wavepackets (when used as a basis) could be applied, for example, to represent Herzberg–Teller initial states and to compute internal conversion rates.69
@@
Optimized $S_{0}$ and $S_{1}$ geometries of anthracene with computed and experimental frequencies of selected vibrational modes, additional computational details of the empirical Duschinsky coupling and of the local harmonic calculation, computed and experimental $0^{0}$ emission spectra, computed spectra in figs 1 and 4 without wavenumber scaling.

### v5

@@ 5 Conclusions
Because the vibrational Hagedorn dynamics in the global harmonic model is exact, it is easier to identify and isolate errors. Our analysis of the results with different harmonic approximations demonstrates the importance of including both mode distortion and Duschinsky rotation in evaluating molecular vibronic spectra, which serve as a sensitive probe for molecular structure and dynamics. For anthracene, our results suggest that the electronic structure methods we used underestimated mode mixing, which affected the accuracy of combination band intensities.
The Hagedorn approach can be readily extended to capture mild anharmonic effects through the local harmonic approximation, which requires expensive on-the-fly electronic structure evaluations along the trajectory (though the same propagation can again be used to evaluate spectra from all initial excitations).62 Computing vibronic spectra within different global harmonic models can help selecting appropriate ab initio methods and identifying when anharmonicity becomes significant before more sophisticated quantum dynamics methods are applied.
Even within the global harmonic framework, an extension of our method could be useful for other experimental techniques involving vibrationally excited states, such as vibrationally promoted electronic resonance (VIPER) experiments,91, 92, 93 time-resolved photoelectron spectroscopy,94 and fluorescence-encoded infrared (FEIR) spectroscopy.95 More generally, Hagedorn functions form a complete orthonormal basis and can exactly expand arbitrary polynomials times a Gaussian.53, 46 This could enable Hagedorn wavepackets to treat more complex initial states, for example, in non-Condon spectroscopy92, 30, 37, 96. Finally, to account for non-adiabatic effects, one could extend the approach by variationally propagating multiple Hagedorn bases on coupled electronic surfaces.97, 98, 99, 56
Data Availability
In addition to the data provided in the article and the Supporting Information, the electronic structure results used to construct the initial state and the global harmonic models, along with the computed spectra and autocorrelation functions, are available on Zenodo (DOI: 10.5281/zenodo.15693437).
@@
Expressions to compute the overlaps between Hagedorn functions, optimized $S_{0}$ and $S_{1}$ geometries of anthracene with computed and experimental frequencies of selected vibrational modes, additional computational details of the empirical Duschinsky coupling and of the local harmonic calculation, computed and experimental $0^{0}$ emission spectra, computed spectra in figs 1 and 4 without wavenumber scaling.